A pair of fair coins is thrown repeatedly until a pair of tails is obtained. The random variable X denotes the number of throws required to obtain a pair of tails.
- (a)Find the expected value of X.[1]
- (b)Find the probability that exactly 3 throws are required to obtain a pair of tails.[1]
- (c)Find the probability that fewer than 6 throws are required to obtain a pair of tails.[2]
- (d)On a different occasion, a pair of fair coins is thrown 80 times. Use an approximation to find the probability that a pair of tails is obtained more than 25 times.[5]
Mark scheme
Marking points from the official mark scheme (May/June 2020). The mark or mark type is shown in brackets.
- [B1] 1/(1/4) = 4
- [B1] (3/4)^2 * (1/4) (=0.141)
- [M1] P(X<6)=1-(3/4)^5
- [A1] 0.763
- [M1] Mean = 80×0.25= 20
- [M1] Var = 80×0.25×0.75 = 15
- [M1] P(more than 25) = P(z > (25.5-20)/√15)
- [M1] P(z > 1.42)
- [M1] 1-0.9222
- [A1] 0.0778
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →